### v1

# Collaborative and parametric insurance on the Ethereum blockchain

Abstract
This paper introduces a blockchain-based insurance scheme that integrates parametric and collaborative elements. A pool of investors, referred to as surplus providers, locks funds in a smart contract, enabling blockchain users to underwrite parametric insurance contracts. These contracts automatically trigger compensation when predefined conditions are met. The collaborative aspect is embodied in the generation of tokens, which are distributed to both surplus providers and policyholders. These tokens represent each participant’s share of the surplus and grant voting rights for management decisions. The smart contract is developed in Solidity, a high-level programming language for the Ethereum blockchain, and deployed on the Sepolia testnet, with data processing and analysis conducted using Python. In addition, open-source code is provided and main research challenges are identified, so that further research can be carried out to overcome limitations of this first proof of concept.
MSC 2010: 91B30.
Keywords: Blockchain, parametric insurance, actuarial science.
1 Introduction
A blockchain is a decentralized ledger maintained by consensus within a peer-to-peer network. The nodes in this network agree on a shared history of data by executing an algorithm known as a consensus protocol. The Bitcoin blockchain, as described by Nakamoto [22], employs an algorithm called proof-of-work (PoW) to validate and record transactions between users. In contrast, the Ethereum blockchain, introduced by Buterin [5], uses a proof-of-stake (PoS) mechanism. Ethereum is designed to operate as a "world computer," enabling developers to deploy compiled code to power decentralized applications (dApps) that users can interact with directly.
Through these decentralized applications, Ethereum extends the capabilities of blockchain beyond simple fund transfers, enabling more complex financial operations and fostering the emergence of a new financial ecosystem known as decentralized finance (DeFi), as described by Schär [24]. Traditional finance relies on trusted third parties and centralized databases, which involve intermediary costs, lack transparency, operate only during limited hours, and are susceptible to cyberattacks. In contrast, distributed ledgers are publicly accessible, replicated in multiple locations, and function continuously, 24/7. Furthermore, the code that governs decentralized applications is auditable, improving transparency. For further discussion, we refer the reader to Lipton and Treccani [19, Chapter 2].
Some criticisms of the financial system can also be extended to the insurance sector. Insurance policies are often complex and difficult for the general public to understand. Premium calculations are typically based on historical data that are privately held by insurance companies, which limits transparency. In addition, the claim settlement process is frequently lengthy and requires on-site assessments by one or more experts. In response to these challenges, the industry has witnessed the emergence of a new type of insurance known as parametric or index-based insurance. For discussion of its application to crop insurance, see Abdi et al. [1] and Conradt et al. [6]. Parametric insurance provides compensation based on predefined triggers—such as specific weather conditions or flight delays—rather than relying on traditional, often time-consuming, claims assessments. Policyholders may, however, be exposed to basis risk—the risk that the index-based compensation does not accurately reflect the actual loss experienced by the policyholder. This straightforward mechanism can be easily encoded in smart contracts, making parametric insurance a promising candidate for blockchain-based insurance solutions.
We present a smart contract on the Ethereum blockchain to manage a portfolio of parametric insurance policies. Our framework is inspired by the discussion in Cousaert et al. [7], from which we adopt the role distribution, while introducing an additional mechanism for the "tokenization" of risk. The smart contract owner writes the code and sets the initial values for parameters such as the premium loading and the risk level required for risk capital allocation. Investors participate by locking funds in the smart contract, in exchange for which they receive protocol tokens. These funds are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. The contribution of each investor is reflected in their token holdings.
Future policyholders interact with the protocol to seek protection against unfortunate events. An insurance agreement can only be concluded if the funds provided by the investors are sufficient to cover the risk. The smart contract funds must remain above a specified threshold at all times, as prescribed by standard insurance regulations such as the European directive Solvency II. A portion of the premiums paid by customers is used to reward them with protocol tokens, which can be exchanged for ETH at any time. The value of the token, or equivalently its exchange rate against ETH, fluctuates over time depending on whether compensation is paid.
The token ownership of policyholders incentivizes honest behavior. Token issuance through risk underwriting decreases the token’s value, encouraging investors to remain committed until certain insurance policies are resolved. Conversely, the token value increases if no compensation is paid at settlement. Additionally, token ownership enables governance of the insurance protocol. The contract owner may update the smart contract parameters based on recommendations from a board composed of token holders, including both investors and policyholders. This governance model transforms the smart contract into a Decentralized Autonomous Organization (DAO).
The tokenization of risk on the blockchain can be viewed as a novel way to design insurance-linked securities, as outlined in Barrieu and Albertini [4]. By enabling policyholders to become investors, our smart contract creates a collaborative insurance scheme, akin to those described in Feng [12]. Investors, referred to as Surplus Providers (SPs), are analogous to liquidity providers in decentralized exchanges, which serve as blockchain-based trading venues for crypto assets; see, for instance, Mohan [21].
We showcase the concrete implementation of our framework using the Solidity programming language. This includes instructions on deploying the contract on the Ethereum testnet, interacting with it, and retrieving data via the Etherscan Application Programming Interface (API). For a comprehensive overview of how the Ethereum blockchain operates and a tutorial on coding in Solidity, we refer the reader to the book of Antonopoulos and Gavin Wood Ph. [2].
We aim to design our smart contract to operate with minimal human interference, maximizing automation. Every interaction with the smart contract incurs a cost, as any transaction submitted to the blockchain requires a fee to be processed by the network of nodes. Therefore, optimizing the number of transactions is crucial for reducing operational costs.
Actuarial methods for calculating risk capital are often too complex to be directly implemented in a high-level language like Solidity, which, for instance, only supports integer-based arithmetic. Additionally, deploying a smart contract on the blockchain entails costs that grow with its complexity, emphasizing the need for simplicity in design. One of our key contributions is the simplification of these computations through justified approximations. In particular, we propose a straightforward recursive formula to determine the solvency capital required to adequately cover the risks of a parametric insurance portfolio. Another contribution of this paper is the demonstration of the feasibility of deploying a beta version of such a smart contract, where certain complexities are intentionally simplified. We also identify the main limitations of this simplified framework and provide the research and actuarial communities with an open-source beta version of the smart contract. This prototype serves as a foundation for developing generalizations and improved versions that address these limitations.
A smart contract is a system defined by state variables that evolve based on actions performed by users. Section 2 introduces the roles of the participants, including the contract owner, surplus providers, and policyholders. It also establishes a probabilistic framework for the state variables using stochastic processes. Examples of state variables include the smart contract’s balance and the tokens held by each participant. Users interact with the smart contract by invoking functions, and the possible actions and their impacts on the system’s state are detailed in Section 3. Section 4 defines the risk capital levels that the surplus should not fall below. It also outlines a method to compute the solvency capital required dynamically as new insurance contracts are added to the portfolio. Section 5 demonstrates how the smart contract operates through an event-driven scenario, showcasing all its functions and culminating in the liquidation and reset of the smart contract. The main limitations of this beta version of the smart contract are discussed in Section 6, along with potential ideas for addressing them.
2 Parametric insurance portfolio management framework
A smart contract defines a system whose state evolves based on the actions of blockchain users. This section describes the roles of the users of our smart contract in Section 2.1, followed by the definition of stochastic processes used to monitor the state of the system in Section 2.2.
2.1 Role distribution
The key participants in our blockchain insurance scheme are the smart contract owner, the investors also referred to as surplus providers, and the policyholders. Both the underwriting and claim assessment processes are fully automated within the smart contract protocol. Consequently, unlike in Cousaert et al. [7], our framework does not include claim assessors or underwriters. It is also worth noting that the role of our investors closely resembles that of reinsurers in Cousaert et al. [7].
2.1.1 Smart contract owner
The smart contract owner is responsible for designing the insurance protocol by writing the Solidity code for the smart contract. After deploying the smart contract on the Ethereum blockchain, the owner initializes two key parameters: the premium loading and the risk level associated with the solvency capital. The premium loading is divided into two components. The first component covers the management costs of the smart contract and provides returns for the investors. The second component is redistributed to the policyholders in the form of protocol tokens.
2.1.2 Surplus providers
Surplus providers, or investors, contribute financially to the smart contract to build solvency capital. These funds are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. In exchange for making their funds available, investors receive tokens that represent their share of the surplus. Surplus providers can withdraw a portion of their funds by burning tokens in exchange for an equivalent amount of ETH, calculated based on the current exchange rate. This exchange rate evolves over time, influenced by whether claims are being compensated. Withdrawals are permitted as long as the funds locked in the smart contract do not fall below a threshold known as the Solvency Capital Requirement (SCR), which corresponds to a quantile of the future loss distribution.
2.1.3 Policyholders
Policyholders seek protection against unfortunate events. If such an event occurs, the smart contract pays a lump-sum compensation denominated in ETH. To obtain this protection, policyholders pay a premium to the smart contract, also denominated in ETH. A small portion of this premium is redistributed to them in the form of protocol tokens. An insurance agreement can only be concluded if the value of the Solvency Capital Requirement (SCR), after including the new risk, does not exceed the current funds locked in the smart contract.
[…]
2.1.4 Decentralized Autonomous Organization
Token ownership grants the right to participate in the protocol’s governance. Token holders collectively form a board that can issue recommendations to modify the protocol, such as adjusting parameters like the safety loading or the level of risk. The responsibility for implementing these changes lies with the smart contract manager, who may also be one of the investors.
2.2 Mathematical framework
In this section, we introduce several stochastic processes to monitor the state of our smart contract. We first focus on the stochastic processes relevant to the management of the parametric insurance portfolio in Section 2.2.1. Then, we examine the stochastic processes associated with the balance of the smart contract and protocol token ownership in Section 2.2.2.
2.2.1 Parametric insurance portfolio over time
A parametric insurance contract $(A_{T},l)$ pays a predefined compensation $l$ if a specific event $A_{T}$ occurs at time $T>0$. Denote by $S$ the time at which the contract is underwritten. Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a filtered probability space such that the event $A_{T}$ is $\mathcal{F}_{T}$-measurable. Define the status of the contract over time as:
$\xi_{t}=\begin{cases}0,&\text{if }t<T,\\ 1,&\text{if }t\geq T\text{ and }\bar{A}_{T},\\ 2,&\text{if }t\geq T\text{ and }A_{T},\\ \end{cases}$ (1)
where the state space of $(\xi_{t})_{t\geq 0}$ is $\mathcal{S}=\{0,1,2\}$. Each state corresponds to a specific condition of the contract:
[…]
Remark 2.1.
The smart contract may be forced to file for bankruptcy if the surplus falls below the minimum capital requirement at any point between $S$ and $T$. If this occurs, the policy is canceled, and the premium is refunded to the policyholder. In such cases, a fourth possible state could be introduced:
$3=\text{"canceled"}.$" .
However, to simplify the presentation, we do not consider this possibility at this stage.
Let $\theta=\mathbb{P}(A_{T})$ denote the probability of the event $A_{T}$. The pure premium associated with a contract $(A_{T},l)$ is given by
$\mu=\theta\cdot l,$
where $\theta=\mathbb{P}(A_{T})$. Let $S<T$ the time at which the policy is underwritten. The commercial premium is then given by
$\pi=(1+\eta_{S})\cdot\mu,$
using the expectation premium principle, where $\eta_{S}>0$ is the loading factor at time $S$ to ensure profitability on average and cover management costs. This loading is split into two components, $\eta_{S}=\eta_{S}^{(1)}+\eta_{S}^{(2)}$, where $\eta_{S}^{(1)}$ covers the protocol’s operational costs and rewards the investors, while $\eta_{S}^{(2)}$ is redistributed to the policyholder in the form of protocol tokens according to an exchange rate defined in Section 2.2.2. The safety loading may evolve over time depending on the decisions of the board of token owners. The processes $\left(\eta^{(1)}_{t}\right)_{t\geq 0}$, $\left(\eta^{(2)}_{t}\right)_{t\geq 0}$ and $\eta_{t}=\eta_{t}^{(1)}+\eta_{t}^{(2)},\text{ }t\geq 0$ are $\mathcal{F}_{t}$-adapted stochastic processes.
A parametric insurance portfolio is a collection of parametric insurance contracts $\{(A_{T_{i}},l_{i})\}_{i\geq 1}$ defined on the probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$. Let $S_{i}$ denote the time at which contract $i$ is underwritten. The number of contracts underwritten up to time $t\geq 0$ is modeled as an $\mathcal{F}_{t}$-adapted counting process, defined as
$N_{0}=0,\quad N_{t}=\sum_{i\geq 1}\mathbb{I}_{S_{i}\leq t}.$
The statuses of the contracts are represented by the vector $\Xi_{t}=\left(\xi_{t}^{(1)},\ldots,\xi_{t}^{(N_{t})}\right)$. Denote by $Z_{i}=\mathbb{I}_{A_{T_{i}}}l_{i}$ the payout of insurance contract $i$, where $(\mathbb{I}_{A_{T_{i}}})_{i=1,\ldots,N_{t}}$ is a sequence of Bernoulli variables, $\text{Bernoulli}(\theta_{i})$, for $i=1,\ldots,N_{t}$.
The liability at time $t>0$ is a stochastic process defined as
$L_{0}=0,\quad L_{t}=\sum_{i=1}^{N_{t}}Z_{i}\mathbb{I}_{S_{i}\leq t,T_{i}>t},\quad\text{for }t\geq 0,$ (2)
it corresponds to the future losses given the open contracts in our portfolio at time $t$. The premiums collected to compensate for this liability are given by
$\Pi_{t}=\sum_{i=1}^{N_{t}}\pi_{i}\mathbb{I}_{S_{i}\leq t,T_{i}>t},\quad\text{for }t\geq 0,$
where
$\pi_{i}=(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i},\quad\text{for }i=1,\ldots,n.$
We refer to $\Pi_{t}$ as the collected but yet to be earned premiums. We define two solvency capitals as
$SCR_{t}=\text{Quantile}(L_{t}-\Pi_{t};\alpha_{\text{SCR}}),\quad\text{ and }MCR_{t}=\text{Quantile}(L_{t}-\Pi_{t};\alpha_{\text{MCR}}).$
Here, $SCR$ stands for a simplified version of Solvency Capital Requirement, and $MCR$ stands for Minimum Capital Requirement, by analogy with the European directive Solvency II, under which $\alpha_{\text{SCR}}=0.995$ and $\alpha_{\text{MCR}}=0.85$.
2.2.2 Balance of the smart contract and token holdings
When deployed on the Ethereum blockchain, a smart contract is assigned an Ethereum address, enabling it to receive and send ETH. The balance $(B_{t})_{t\geq 0}$ increases each time an investor contributes funds to the contract or when a policy is underwritten and a premium is paid. Conversely, the balance decreases when an investor withdraws funds or when a compensation is paid to a policyholder in the event of a covered loss. The balance decreases when an investor withdraws some funds from the contract and when a compensation is paid to an unfortunate policyholder. Let $(\tau_{i}^{+})_{i\geq 1}$ and $(\tau_{i}^{-})_{i\geq 1}$ be the times at which investors make deposits and withdrawals, respectively. Denote the associated amounts by $(x_{i}^{+})_{i\geq 1}$ for deposits and $(x_{i}^{-})_{i\geq 1}$ for withdrawals. We have
$B_{0}=0,\text{ and }B_{t}=\sum_{i\geq 1}\mathbb{I}_{\tau_{i}^{+}\leq t}x_{i}^{+}-\sum_{i\geq 1}\mathbb{I}_{\tau_{i}^{-}\leq t}x_{i}^{-}+\sum_{i\geq 1}\mathbb{I}_{S_{i}\leq t}\pi_{i}-\sum_{i\geq 1}\mathbb{I}_{T_{i}\geq t}\mathbb{I}_{A_{T_{i}}}l_{i},\text{ for }t\geq 0.$
We define $(X_{t})_{t\geq 0}$ as the surplus of the smart contract, representing the funds available to underwrite new insurance policies. Unlike the balance of the smart contract, the surplus does not increase when a new contract is underwritten. Instead, the premium is added at the time of the policy’s resolution, when it is considered "earned" rather than merely "collected". Therefore we have
$X_{0}=0,\text{ and }X_{t}=\sum_{i\geq 1}\mathbb{I}_{\tau_{i}^{+}\leq t}x_{i}^{+}-\sum_{i\geq 1}\mathbb{I}_{\tau_{i}^{-}\leq t}x_{i}^{-}+\sum_{i\geq 1}\mathbb{I}_{T_{i}\leq t}\pi_{i}-\sum_{i\geq 1}\mathbb{I}_{T_{i}\geq t}\mathbb{I}_{A_{T_{i}}}l_{i}\text{ for }t\geq 0.$
Note that $B_{t}\geq X_{t}$ for every $t\geq 0$. The level of $(X_{t})_{t\geq 0}$, representing the surplus of the smart contract, is compared to $SCR_{t}$ (Solvency Capital Requirement) and $MCR_{t}$ (Minimum Capital Requirement). If $X_{t}\leq SCR_{t}$, no new policies may be underwritten, and investors are prohibited from withdrawing funds. If $X_{t}\leq MCR_{t}$, the smart contract ceases its operations, and the remaining balance is redistributed—first to the policyholders and then to the investors in a way detailed in Section 3. The active insurance policy at time of termination of the smart contract are being canceled as mentioned in Remark 2.1.
Whenever the smart contract receives funds from external sources, protocol tokens are issued. Let $(M_{t})_{t\geq 0}$ denote the total number of distinct Etehereum addresses that have ever transferred funds to the smart contract, whether as investments or premium payments. Let
$Y_{t}=\left(\begin{array}[]{ccc}Y_{t}^{(1)}&\ldots&Y_{t}^{(M_{t})}\end{array}\right),$
be a vector representing the token holdings of the smart contract participants. The total supply of tokens is then given by $\bar{Y}_{t}=\sum_{i=1}^{M_{t}}Y_{t}^{(i)}$. The exchange rate of tokens against ETH is modeled by a stochastic process $(r_{t})_{t\geq 0}$, defined as
$r_{0}=1,\quad r_{t}=\frac{X_{t}}{\bar{Y}_{t}},\quad\text{for }t\geq 0.$
The exchange rate must be updated every time $X_{t}$ or $\bar{Y}_{t}$ changes. If, at time $t$, an amount $x$ of funds is transferred to the smart contract, a corresponding quantity of tokens $y=x/r_{t}$ is issued. Conversely, if a quantity $y$ of tokens is burned at time $t>0$, the smart contract transfers $x=y\cdot r_{t}$ back to the transaction issuer.
Section 3 provides a detailed account of the dynamics of the various stochastic processes we have just introduced.
3 A state description of the smart contract
A smart contract defines a system whose state evolves based on the actions of blockchain users. In our parametric insurance example, these users include the contract owner, the surplus providers, and the policyholders (see Section 2.1). Participants interact with the smart contract through functions that modify the state of the system. The remainder of this section describes each possible action and its impact on the stochastic processes introduced in Section 2.2.
3.1 Funding the contract
Assume that an investor transfers $x$ ETH to the contract at time $t>0$. Both the balance of the contract and the surplus increase as follows:
$B_{t+h}=B_{t}+x,\quad X_{t+h}=X_{t}+x,$
where $h>0$ is an infinitesimal time lapse. If the investor is new, then $M_{t+h}=M_{t}+1$; otherwise, $M_{t+h}=M_{t}$. The $j$-th component ($j\in\{1,\ldots,M_{t+h}\}$) of the vector $(Y_{t})_{t\geq 0}$ is updated as:
$Y_{t+h}^{(j)}=Y_{t}^{(j)}+y,$
where $y=x/r_{t}$.
3.2 Underwrite a parametric insurance policy
A customer wishes to purchase an insurance contract $(A_{T},l)$ at time $t=S$. The premium for such a parametric insurance contract is given by:
$\pi=(1+\eta_{t})\cdot l\cdot\theta,$
where $\theta=\mathbb{P}(A_{T})$. The contract can only be underwritten if the surplus of the smart contract is sufficient to cover the resulting increase in the solvency capital requirement. The liability and the cumulative premium collected would then be updated to:
$\widetilde{L}_{t}=L_{t}+l\cdot\mathbb{I}_{A_{T}},\quad\widetilde{\Pi}_{t+h}=\Pi_{t}+\pi.$
We calculate the updated solvency capital requirement:
$\widetilde{\text{SCR}}_{t}=\text{Quantile}(\widetilde{L}_{t}-\widetilde{\Pi}_{t};\alpha),$
and underwrite the contract only if:
$X_{t}\geq\widetilde{\text{SCR}}_{t}.$ (3)
If condition (3) holds, the number of contracts, liability, and premium collected are updated as follows:
$N_{t+h}=N_{t}+1,\quad L_{t+h}=L_{t}+l_{N_{t+h}}\cdot\mathbb{I}_{A_{T_{N_{t+h}}}},\quad\Pi_{t+h}=\Pi_{t}+(1+\eta_{t})\cdot l_{N_{t+h}}\cdot\theta_{N_{t+h}}.$
The SCR and MCR are also updated as:
$\text{SCR}_{t+h}=\text{Quantile}(L_{t+h}-\Pi_{t+h};\,\alpha_{\text{SCR}}),\quad\text{MCR}_{t+h}=\text{Quantile}(L_{t+h}-\Pi_{t+h};\,\alpha_{\text{MCR}}),$
using the recursive method based on a normal approximation of the distribution of $L_{t}$ described in Section 4.
We add a component:
$\xi_{t+h}^{(N_{t+h})}=0,$
to the vector $\Xi_{t}$ that gathers the statuses of the insurance policies. The balance of the contract is updated as
$B_{t+h}=B_{t}+(1+\eta_{t})\cdot\theta\cdot l,$
while the surplus $X_{t}$ remains unchanged. The policyholder receives:
$y=\frac{\eta_{t}^{(2)}\cdot\theta\cdot l}{r_{t}},$
in protocol tokens. If the policyholder is neither an investor nor a former policyholder, the number of token holders increases as
$M_{t+h}=M_{t}+1,$
otherwise, $M_{t+h}=M_{t}$. A component $j\in\{1,\ldots,M_{t+h}\}$ of the vector $Y_{t}$ is updated as
$Y_{t+h}^{(j)}=Y_{t}^{(j)}+y.$
Finally, the exchange rate becomes
$r_{t+h}=\frac{X_{t+h}}{\bar{Y}_{t+h}}=\frac{X_{t}}{\bar{Y}_{t+h}+y}.$
3.3 Withdrawing from the smart contract and burning tokens
An investor or policyholder $j\in\{1,\ldots,M_{t}\}$ may withdraw some of their funds by burning protocol tokens at time $t>0$. This action is subject to the conditions that the participant holds enough protocol tokens $Y^{(j)}_{t}$, the smart contract has a sufficient balance $B_{t}$, and the surplus $X_{t}$ remains adequate. Assume that participant $j$ wishes to burn $y$ tokens at time $t>0$. Upon burning, the participant should receive $x=y\cdot r_{t}$ ETH. The transaction is allowed if
$y\leq Y^{(j)}_{t}\text{, }x\leq B_{t},\text{ and }x<X_{t}-SCR_{t}.$ (4)
Provided that (4) is satisfied then we update the contract balance and the surplus as
$B_{t+h}=B_{t}-x,\text{ and }X_{t+h}=X_{t}-x.$
We further update the token balance of investor $j$ with
$Y_{t+h}^{(j)}=Y_{t}^{(j)}-y.$
3.4 Contract resolution and claim settlement
At time $t=T_{i}$, contract $i$, for some $i=1,\ldots,N_{t}$, is settled as it becomes known whether the event $A_{T_{i}}$ has occurred.
Let us first consider the simple case where $A_{T_{i}}$ does not occur. In this scenario, the number of contracts, liability, and premiums collected are updated as follows,
$N_{t+h}=N_{t}-1,\quad L_{t+h}=L_{t}-l_{i},\quad\Pi_{t+h}=\Pi_{t}-(1+\eta_{S_{i}})\cdot l_{i}\cdot\theta_{i}.$
Recall that $S_{i}$ is the time at which contract $i$ was underwritten. The SCR and MCR are also updated as
$\text{SCR}_{t+h}=\text{Quantile}(L_{t+h}-\Pi_{t+h};\alpha_{\text{SCR}}),\quad\text{MCR}_{t+h}=\text{Quantile}(L_{t+h}-\Pi_{t+h};\alpha_{\text{MCR}}),$
using the recursive method based on a normal approximation of the distribution of $L_{t}$ described in Section 4. The balance of the contract remains unchanged $B_{t+h}=B_{t}$, while the surplus is updated as
$X_{t+h}=X_{t}+(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i}.$
The premium for contract $i$ is now considered "earned." The exchange rate is updated as
$r_{t+h}=\frac{X_{t+h}}{\bar{Y}_{t+h}},$
where it is noted that the total supply of protocol tokens remains unchanged between $t$ and $t+h$.
If $A_{T_{i}}$ occurred, the balance, number of contracts, liability, and premiums collected are updated as follows,
$\widetilde{N}_{t}=N_{t}-1,\quad\widetilde{L}_{t}=L_{t}-l_{i},\quad\widetilde{\Pi}_{t}=\Pi_{t}-(1+\eta_{S_{i}})\cdot l_{i}\cdot\theta_{i}.$
The SCR and MCR are also updated as
$\widetilde{\text{SCR}}_{t}=\text{Quantile}(\widetilde{L}_{t}-\widetilde{\Pi}_{t};\alpha_{\text{SCR}}),\quad\widetilde{\text{MCR}}_{t}=\text{Quantile}(\widetilde{L}_{t}-\widetilde{\Pi}_{t};\alpha_{\text{MCR}}).$
of the contract and the surplus are updated as
$\widetilde{B}_{t}=B_{t}-l_{i},\text{ and }\widetilde{X}_{t}=X_{t}-l_{i}+(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i}.$
The use of the tilde indicates that this update is not necessarily final, as the surplus may cross the MCR threshold
$\widetilde{X}_{t}\leq\text{MCR}_{t+h}.$
If this occurs, the smart contract is reset. The outstanding balance is first used to reimburse the policyholders of ongoing contracts. The premium refunded to the policyholder is the amount paid, excluding the participation loading that granted tokens to the policyholder. The total amounts is therefore given by
$x=\sum_{i=1}^{\widetilde{N}_{t}}\mathbb{I}_{\xi^{(i)}_{t}=0}\left(1+\eta_{S_{i}}^{(1)}\right)\cdot\theta_{i}\cdot l_{i}.$
If $\widetilde{X}_{t}>x$, the premiums are fully refunded, and the remaining wealth, $\widetilde{X}_{t}-x$, is distributed to the investors according to their token holdings, using the updated exchange rate
$\widetilde{r}_{t}=\frac{\widetilde{X}_{t}-x}{\bar{Y}_{t}}.$
The smart contract transfers
$x_{j}=\widetilde{r}_{t}\cdot Y^{(j)}_{t},$
to all investors $j=1,\ldots,M_{t}$.
If $\widetilde{X}_{t}\leq x$, the policyholders are reimbursed pro rata based on the premiums they paid. Specifically, we define weights for the policyholders with active contracts as
$w_{i}=\frac{\mathbb{I}_{\xi^{(i)}_{t}=0}\left(1+\eta_{S_{i}}^{(1)}\right)\cdot\theta_{i}\cdot l_{i}}{\sum_{i=1}^{\widetilde{N}_{t}}\mathbb{I}_{\xi^{(i)}_{t}=0}\left(1+\eta_{S_{i}}^{(1)}\right)\cdot\theta_{i}\cdot l_{i}},$
and transfer
$x_{i}=w_{i}\cdot\widetilde{X}_{t},$
to policyholder $i$ for $i=1,\ldots\widetilde{N}_{t}$.
After completing these transfers, some smart contract variables are reset to their initial state with
$B_{t+h}=0,\quad X_{t+h}=0,\quad L_{t+h}=0,\quad\Pi_{t+h}=0,\quad\text{SCR}_{t+h}=0\quad\text{MCR}_{t+h}=0\quad r_{t+h}=1.$
The number of contracts and the number of token holders remain unchanged:
$N_{t+h}=N_{t},\quad M_{t+h}=M_{t}.$
However, all token holdings are reset to zero:
$Y^{(j)}_{t+h}=0,\quad\text{for all }j=1,\ldots,M_{t+h}.$
The practical implementation of the smart contract using the programming language solidity is documented in the online accompaniement of this paper [^3] https://github.com/LaGauffre/smart_parametric_insurance/blob/main/latex/sup_material_blockchain_parametric_insurance.pdf.
4 Recursive calculation of the solvency capitals
Recall that the liability at time $t>0$ is given by
$L_{0}=0,\quad L_{t}=\sum_{i=1}^{N_{t}}Z_{i}\mathbb{I}_{S_{i}\leq t,T_{i}>t},\quad\text{for }t\geq 0,$ (5)
where $Z_{i}=l_{i}\mathbb{I}_{A_{T_{i}}}$ for $i\geq 0$. Assume that the Bernouilli variables $\left(\mathbb{I}_{A_{T_{i}}}\right)_{i\geq 0}$ are mutually independent. Then, provided that $\sum_{i=1}^{N_{t}}\mathbb{I}_{S_{i}\leq t,T_{i}>t}=n$, the distribution of $L_{t}$ corresponds to that of the aggregate losses in the individual risk model described in standard actuarial textbooks such as Klugman [18, Chapter 9, Section 11]. While the exact distribution may be computed via combinatorial analysis, this approach becomes impractical as $n$ grows large. A more efficient method is to use a recursive algorithm like that of De Pril [8]. 1 recalls this algorithm adapted to our notations.
Proposition 1.
Assume that $l_{i}\in\mathbb{N}$ for $i=1,\ldots,n$ then we have
$\mathbb{P}(L_{t}=0)=\prod_{i=1}^{n}(1-p_{i}),$
and
$\mathbb{P}(L_{t}=x)=\sum_{i=1}^{n}v_{i}(x),$
where
$v_{i}(x)=\begin{cases}0,&x<l_{i},\\ \frac{p_{i}}{1-p_{i}}l_{i}\left[\mathbb{P}(L_{t}=x-l)-v_{i}(x-l_{i})\right],&\text{ otherwise,}\end{cases}$
for $i=1,\ldots,n$.
Once the probability distribution of $L_{t}$ is obtained, we can compute its cumulative distribution function and use a bisection algorithm to determine the quantile of interest. While this method is significantly more efficient than combinatorial analysis, it can still be time-consuming for large $n$ and large values of $l_{i}$. Moreover, numerical instabilities are common in such recursive formulas, often arising from operations on very small probabilities or the subtraction of nearly equal values. In such cases, one can consider approximation methods, such as the approach proposed by Hipp [17], which uses a compound Poisson distribution as an approximation. The distribution of the compouund Poisson distribution then requires the use of Panjer’s algorithm of Panjer [23] or a fast Fourier transform inversion method, see Embrechts and Frei [11] for a comparison with Panjer’s recursion.
The numerical procedures mentioned above are effective but not very practical for automated implementation on the blockchain. As noted in the introduction, smart contract functions cannot encode complex operations. This limitation means that solvency capital calculations would need to be performed off-chain, requiring intervention from the contract manager. Moreover, any change in the portfolio at time $s\geq t$ would necessitate recomputation of the probability distribution of the liability $L_{s}$ to determine $\text{SCR}_{s}$ or $\text{MCR}_{s}$.
For instance, if a policy $(A_{T},l)$ is added or removed (e.g., due to settlement, in which case $s=T$), we would need a method to update the SCR efficiently. To achieve this, we rely on a normal approximation, justified by a generalized central limit theorem, which applies to independent random variables with finite variance. Denote by
$\mathbb{E}(Z_{i})=\mu_{i}=\theta_{i}\cdot l_{i},\text{ and }\mathbb{V}(Z_{i})=\sigma_{i}^{2}=\theta_{i}\cdot(1-\theta_{i})\cdot l_{i}^{2}\text{ for }i=1,\ldots,n,$
the mean and variance of the $Z_{i}$’s. Let $s_{n}^{2}=\sum_{i=1}^{n}\sigma_{i}^{2}$ be the sum of the variances. Suppose that the sequence of random variables $(Z_{i})_{i=1,\ldots,n}$ satisfies
$\underset{n\rightarrow\infty}{\lim}\frac{1}{s_{n}^{2}}\sum_{i=1}^{n}\mathbb{E}\left[\left(Z_{i}-\mu_{i}\right)^{2}\mathbb{I}_{|Z_{i}-\mu_{i}|>\epsilon s_{n}^{2}}\right]=0,\text{ for every }\epsilon>0,$ (6)
then we have
$\frac{1}{s_{n}}\sum_{i=1}^{n}(Z_{i}-\mu_{i})\overset{\mathcal{D}}{\longrightarrow}\mathsf{Normal}(0,1),\text{ as }n\rightarrow\infty,$
where $\mathsf{Normal}(0,1)$ is the standard normal distribution and $\overset{\mathcal{D}}{\longrightarrow}$ means convergence in distribution.
Proposition 2.
Suppose that the $l_{i}$’s are bounded, then condition 6 holds.
Proof.
The Lyapunov condition holds if there exists $\delta>0$ such that
$\underset{n\rightarrow\infty}{\lim}\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left(|Z_{i}-\mu_{i}|^{2+\delta}\right)=0,$ (7)
and implies condition (6). Denote by $l=\sup_{i\geq 0}l_{i}$ then we have
$\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left(|\xi_{i}-\mu_{i}|^{2+\delta}\right)$$\leq$ $\frac{l^{\delta}}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left(|\xi_{i}-\mu_{i}|^{2}\right)$
$=$ $\left(\frac{l}{s_{n}}\right)^{\delta}\rightarrow 0\text{ for every }\delta>0.$
We note that
$L_{t}-\Pi_{t}=\sum_{i}^{n}Z_{i}-\sum_{i=1}^{n}(1+\eta_{S_{i}})\cdot\mu_{i}=s_{n}\left(\frac{1}{s_{n}}\sum_{i}^{n}(Z_{i}-\mu_{i})\right)-\sum_{i=1}^{n}\eta_{S_{i}}\cdot\mu_{i}.$
We deduce the following approximations of the SCR and MCR,
$\text{SCR}_{t}\approx s_{n}\cdot q_{\alpha_{\text{SCR}}}-\eta_{S_{i}}\cdot\sum_{i=1}^{n}\mu_{i},\text{ and }\text{MCR}_{t}\approx s_{n}\cdot q_{\alpha_{\text{MCR}}}-\eta_{S_{i}}\cdot\sum_{i=1}^{n}\mu_{i},$
where $q_{\alpha}$ is the quantile of order $\alpha$ of the standard normal distribution. Suppose that at time $t=S$, a contract $(A_{T},l)$ such that $\mathbb{P}(A_{T})=\theta$, is underwritten then we update the SCR as
$\text{SCR}_{t+h}=\sqrt{s_{n}^{2}+\theta\cdot(1-\theta)\cdot l^{2}}\cdot q_{\alpha_{\text{SCR}}}-\left(\sum_{i=1}^{n}\eta_{S_{i}}\mu_{i}+\eta_{S}\cdot\theta\cdot l\right).$
Conversely, suppose that at time $t=T_{k}$ contract $k\in\{1,\ldots,N_{t}\}$ is settled then we update the SCR as
$\text{SCR}_{t+h}=\sqrt{s_{n}^{2}-\theta_{k}\cdot(1-\theta_{k})\cdot l_{k}^{2}}\cdot q_{\alpha_{\text{SCR}}}-\left(\sum_{i=1}^{n}\eta_{S_{i}}\mu_{i}-\eta_{S_{k}}\cdot\theta_{k}\cdot l_{k}\right).$
Note that updating the MCR is performed in the same manner, with $q_{\alpha_{\text{SCR}}}$ replaced by $q_{\alpha_{\text{MCR}}}$. The updates to the SCR and MCR can be implemented recursively in Solidity, as they only require updating $s_{n}^{2}$ and $\sum_{i}\eta_{S_{i}}\mu_{i}$. To achieve this, we implemented a custom square root function in Solidity; details can be found in the online supplementary materials [^4] https://github.com/LaGauffre/smart_parametric_insurance/.
The quantiles of the normal distribution used to determine the SCR and MCR are hardcoded but can be modified by the contract owner upon recommendation from the board of token holders.
Remark 4.1.
We have not addressed the estimation of $\theta$ in detail. Standard actuarial practice would recommend using historical data and a regression approach to account for the characteristics of the insured event. Logistic regression is particularly suitable for this task, especially if the probability of the event’s occurrence is influenced by covariates. This is often the case for weather-related events, as discussed in Figueiredo et al. [13]. The formula for calculating $\theta$ could be hardcoded into the smart contract, with covariate values retrieved via relevant APIs, commonly referred to as data oracles in the blockchain ecosystem. Additionally, a function could be included to allow the contract owner to update the regression coefficients if needed.
In cases where the smart contract only underwrites the same type of risk—meaning the Bernoulli variables are independent and identically distributed—a simple beta-binomial credibility model can be used for experience rating; see the monograph of Hans Bühlmann [16, Section 2.1.2] for details.
In our implementation, however, we do not query any API nor include a pricing formula within the smart contract. Instead, we emulate this by generating a random number between 00 and $1$ based on an event description represented as a string. We apply the keccak256 hash function to the string to produce a 256-bit number, which is then converted into a binary sequence of length 256. The resulting number is taken modulo $10{,}000$, and the outcome is divided by $10{,}000$, yielding a pseudo-random probability drawn from a standard uniform distribution.
To determine whether compensation is paid, we process the description of the observed event in the same manner. If the resulting number for the observed event is smaller than that of the original event, a compensation payment is triggered. The getRandomNumber function is provided in the online accompaniement.
A brief illustration of the performance of our normal approximation for estimating the MCR and SCR is provided in 1.
Example 1.
Assume that a collection of contracts $\{(A_{T_{i}},l_{i})\}_{i=1,\ldots,n}$ are underwritten sequentially at times $S_{1}\leq S_{2}\leq\ldots\leq S_{n}$ such that $S_{n}\leq\underset{i=1,\ldots,n}{\min}T_{i}$. Suppose that the probability of the events are given by
$\mathbb{P}(A_{T_{i}})=\theta_{i}\sim\mathsf{Unif}(0,1)\text{ for }i=1,\ldots,n.$
to be consistent with the probability generating process described in Remark 4.1. We also randomly generate the potential compensations with
$l_{i}\sim\mathsf{Unif}(\{1,5\})\text{ for }i=1,\ldots,n,$
[…]
Figure 1: SCR and MCR over time using the normal approximation.
5 Interaction with the smart contract
During its development phase, a smart contract is not immediately deployed on the Ethereum mainnet. Instead, testnets are used to mimic the behavior of the real blockchain, allowing developers to ensure that the smart contract functions as intended. For our project, we deployed the smart contract on the Ethereum testnet called Sepolia. It features a block explorer, known as Etherscan [^5] See https://sepolia.etherscan.io/, which provides an API to retrieve data. Our contract, named ParametricInsurance, is located at the address:
[…]
It can be viewed on Etherscan666see https://sepolia.etherscan.io/address/0xa4f3a8567e1d712fc8a53b3a1b6ec76dd42cc10b. The Solidity code is thoroughly discussed in the online supplementary materials accompanying this paper [^7] https://github.com/LaGauffre/smart_parametric_insurance/.
The deployment of the contract was carried out using Remix IDE [^8] See https://remix.ethereum.org/. Remix IDE is an open-source, web-based integrated development environment (IDE) designed specifically for writing, testing, and deploying smart contracts on the Ethereum blockchain. The Solidity code was first compiled before being deployed to the blockchain. Once deployed, the contract is associated with an Ethereum address, allowing interaction through Remix IDE by calling functions and checking the values of the state variables.
Each action on the testnet incurs a cost denominated in Sepolia testnet tokens, which act as a substitute for ETH. Payments are made through wallets. We use MetaMask [^9] See https://metamask.io/, a popular cryptocurrency wallet and gateway to blockchain applications. MetaMask serves as a browser extension or mobile app that enables users to interact with Ethereum-based decentralized applications (dApps) directly from their web browser or smartphone.
To test the smart contract, three Ethereum accounts were created to represent the roles of the smart contract owner, surplus provider, and policyholder. The addresses of these accounts (plus that of the smart contract) are provided in Table 1.
ETH Address Role Intial balance (in test net ETH)
0xE8e79B8B8c0481fa33a8E0fcA902ad5754BfE1C3 owner $0.4366$
0x2CF8ed1664616483c12Ef3113f6F62E68f1a810A surplus provider $0.2$
0xd34a37613A382bA503f1599F514C9788dF3659C4 policyholder $0.21$
0xa4f3a8567e1d712fc8a53b3a1b6ec76dd42cc10b smart contract 00
Table 1: Ethereum addresses of the smart contract users.
We designed a scenario that calls all the functions available in the smart contract described in Section 3. Here, we take an example involving some weather risks (hail, wind, …) for illustration purposes, but other risks could be considered (flight delays, event cancellation, …). The steps are as follows:
  1. 1.
    Owner deploys the ParametricInsurance contract with the initial parameters:
    $\eta^{(1)}=0.1,\quad\eta^{(2)}=0.05,\quad\alpha_{\text{SCR}}=0.995,\quad\text{and }\alpha_{\text{MCR}}=0.85.$
  2. 2.
    Surplus provider sends ETH 0.1 to the smart contract.
  3. 3.
    Owner sends ETH 0.08 to the smart contract.
  4. 4.
    Policyholder underwrites policy #1 with characteristics ("$rain",0.06\right)$, where:
    $\theta_{1}=\mathbb{P}("rain")=0.7164.$
    Recall from Remark 4.1 that underwriting is done by providing a description of the event as a character string. The function getRandomNumber returns a probability for the event, which is compared to the probability obtained by applying the same function to the description of the observed event.
  5. 5.
    Policyholder underwrites policy #2 with characteristics ("$hail",0.01\right)$, where:
    $\theta_{2}=\mathbb{P}("hail")=0.0315.$
  6. 6.
    Owner updates the parameters:
    $\eta^{(1)}=0.025,\quad\eta^{(2)}=0.025.$
  7. 7.
    Policyholder underwrites policy #3 with characteristics ("$snow",0.1\right)$, where:
    $\theta_{3}=\mathbb{P}("snow")=0.5328.$
  8. 8.
    Owner settles policy #2 with "hail1" as the observed event, where:
    ("$\mathbb{P}("hail1")=0.4496>\theta_{2},$
    resulting in no compensation.
  9. 9.
    Owner burns 0.1 protocol tokens.
  10. 10.
    Owner settles policy #3 with "snow2" as the observed event, where:
    ("$\mathbb{P}("snow2")=0.2058<\theta_{3},$
    triggering a compensation of $l_{3}=\text{ETH}0.1$.
  11. 11.
    Policyholder underwrites policy #4 with characteristics ("$sun",0.08\right)$, where:
    $\theta_{4}=\mathbb{P}("sun")=0.4188.$
  12. 12.
    The Surplus provider burns $0.0157$ protocol tokens, causing the surplus to reach the SCR level. In other words, the Surplus provider withdraws the maximum amount allowable.
  13. 13.
    Owner changes the parameter:
    $\alpha_{\text{MCR}}=0.985.$
  14. 14.
    Owner settles policy #4 with "sun1" as the observed event, where:
    ("$\mathbb{P}("sun1")=0.0128<\theta_{4},$
    triggering a compensation of $l_{4}=\text{ETH}0.08$.
  15. 15.
    Policyholder underwrites policy #5 with characteristics ("$wind",0.02\right)$, where:
    $\theta_{5}=\mathbb{P}("wind")=0.4835.$
  16. 16.
    Owner changes the parameter:
    $\alpha_{\text{MCR}}=0.99.$
  17. 17.
    Owner burns $0.023$ protocol tokens, causing the surplus to reach the SCR level. In other words, the Owner withdraws the maximum amount allowable.
  18. 18.
    Owner settles policy #5 with "wind9" as the observed event, where:
    ("$\mathbb{P}("wind9")=0.1940<\theta_{5},$
    triggering a compensation of $l_{5}=\text{ETH}0.02$.
The final event causes the smart contract to file for bankruptcy, as the surplus $X$ falls below the MCR threshold. Consequently, the premium for policy #$\#1$ is refunded to the policyholder. The remaining surplus is then distributed among all token holders, including the owner, surplus provider, and the policyholder.
We do not claim that the participants in this scenario acted strategically—in fact, quite the opposite. By withdrawing the maximum amount allowable on two occasions, both the Surplus provider and the Owner significantly jeopardized the solvency of the smart contract. Additionally, the parameter adjustments, which made the SCR and MCR nearly identical, are not representative of realistic scenarios. Lastly, the situation where multiple insurance contracts result in compensation is both highly unlucky and statistically improbable. These choices were intentional, designed to illustrate the consequences of the smart contract going bankrupt and undergoing a reset.
Deploying the contract results in a "contract creation" transaction on the blockchain. Each subsequent function call also generates a transaction recorded on the blockchain. These transactions can be retrieved and organized into a data frame, the structure of which is detailed in Table 2.
Variable Name Type Example Value Description
blockNumber int 7143345 Number of the block in the blockchain
blockHash hex 0xf8d9aa8$\ldots$ Hash value of the block
timeStamp int 1732457004 Time at which the transaction got recorded
hash hex 0x93cf64$\ldots$ Hash value of the transaction, servees as identifier on the blockchain
transactionIndex int 10 Location of the transaction in the block
from hex 0xe8e7$\ldots$ Address of the transaction sender
to hex Address of the transaction recipinet, here it is the smart contract address
value int 0 The amount transferred within the transaction
gas int 2935995 gas limit to be spend by the transaction sender
gasPrice int 12145467087 Value of the unit of gas (in WEI where $\text{WEI}1=\text{ETH}10^{-18})$
input data 0x6080604052612$\ldots$ The input data of tthe transaction
methodId hex 0x60806040$\ldots$ Identifier of the function being called
functionName string "Fund" Name of the function being called
contractAddress hex 0xa4f3a856$\ldots$ Address of the smart contract
gasUsed int 2911719 The actual amount of gas used to send the transaction.
Table 2: Description of the transactions data frame associated with the contract.
As mentioned earlier, each transaction incurs a cost to the sender, calculated as:
[…]
The value of a unit of gas (gasPrice) fluctuates over time, as it depends on the level of activity on the Ethereum network. High network congestion leads to higher gas prices. Another factor influencing transaction costs is the amount of data being recorded, which is reflected in the input field of the transaction details.
[…]
Figure 2: Transaction costs in ETH.
[…]
In addition to the list of all transactions, we can retrieve the log of events emitted by the smart contract. A common practice is to associate a specific event with each function. Function calls impact the state of the system, and events provide a mechanism to monitor state transitions effectively. However, care must be taken when defining events, as events that contain a large amount of information will require more data storage and, consequently, increase transaction costs. The definition of the events are provided in Table 3.
Event Name Short Name Logged Information Type Example Value Description
ParametersUpdated Update newEta1 int $1000$ premium loading
newEta2 int $500$
newQAlphaSCR int $25758$ Quantile of the normal distribution
newQAlphaMCR int $10364$
Fund Fund from hex $0x123\ldots$ address of the sender
x int 10000 Amount of ETH
y int 10000 Amount of token
Burn Burn from hex $0x123\ldots$ address of the sender
x int 10000 Amount of ETH
y int 10000 Amount of token
InsuranceUnderwritten Underwrite contractId int 1 Index of the policy
customer hex $0x456\ldots$ Ethereum address
EventDescription string "rain" Description of the insured event
payoutAmount int 1000000 Compensation if the event occurs
status int 0 Status of the insurance policy
ClaimSettled Settle contractId int 1 INdex of the policy
customer hex $0x456\ldots$ Ethereum address
payoutTransferred boolean True Indicates if a compensation has been paid
Xt int 8000000 Surplus
Table 3: Description of Events Generated by the Smart Contract
The events used in this study are defined in the Solidity code, which is included in the online supplementary materials accompanying this article. Using Python, we retrieve the event logs to track participant balances over time. Figure 3 shows the balances (in ETH) of the Ethereum addresses associated with the Owner, the Surplus Provider, the Policyholder, and the Smart Contract after each call to the function.
Figure 3: Balance in ETH of the Ethereum address of the Owner, the Surplus Provider, the policyholder, and the Smart Contract.
We observe that the scenario was particularly beneficial to the Policyholder, who received three compensation payments and a refund for the first policy she purchased. The event logs also enable us to visualize the trajectories of the various processes defined in Section 3. Specifically, Figure 4 displays the trajectories of $X$ (surplus), $B$ (balance), SCR (Solvency Capital Requirement), and MCR (Minimum Capital Requirement) as they evolved throughout the scenario.
Figure 4: Trajectories of $X$, $B$, SCR and MCR.
6 Limitations and perspectives
This article proposes a framework for a simplified smart contract in insurance and demonstrates the feasibility of its implementation on a blockchain. Additionally, it makes this open-source prototype available to other researchers, facilitating further exploration of its limitations and potential enhancements. In this section, we outline the main limitations of this initial approach and suggest possible solutions.
The decentralized nature of the contract and the parametric nature of its trigger offer significant cost-saving opportunities in claim management, distribution, and fraud prevention. For example, in general insurance contracts, fraud—though inherently difficult to quantify—may account for as much as $5\%$ to $15\%$ of collected premiums. In France, insurance fraud was estimated at a minimum of 466 million euros in 2023, representing a $50\%$ increase from 2022101010Source: ALFA (https://www.alfa.asso.fr/), the french Insurance Fraud Control Agency.
Some features that are simplified in this prototype (as complexity increases costs) could be improved without compromising the smart contract’s attractiveness in terms of pricing. However, addressing these improvements must account for potential challenges, such as reputation risks or other factors that might discourage customer participation. Enhancing risk computations, compliance, and risk management must be carefully balanced. Excessive complexity risks becoming counterproductive, akin to the trade-offs observed in the ruin prevention framework discussed by Gauchon et al. [14] and Gauchon et al. [15].
In our simplified approach, claims are assumed to be independent, and the underlying risks are considered stationary. However, in practice, many risks suitable for smart contracts with parametric triggers are influenced by regime changes, which can lead to event clustering or increased claim frequency. To mitigate adverse selection, a critical first step in risk management is introducing a waiting period before the guarantee becomes active. This measure helps prevent policyholders from exploiting arbitrage opportunities when they are aware of worsening risks. For instance, in the case of flight disruptions, events like strikes, conflicts, or volcanic eruptions can make delays almost certain in the short term. Similarly, for weather-related risks, extreme event warnings may be issued days in advance. It is essential to ensure that policyholders cannot exploit the smart contract under such circumstances while paying the same premium as usual.
In some cases, a waiting period may not be sufficient to mitigate risk, as underlying risks can be anticipated several months in advance. For example, the El Ni no/La Ni na cycle, which is partially predictable months ahead, increases the likelihood of certain extreme weather events while reducing others. Moreover, long-term factors such as climate change and sectoral inflation can exacerbate risks over time. If competitors actively adjust their premiums to reflect these changes while the smart contract does not, adverse selection becomes a likely outcome.
To address these challenges, the model should incorporate additional complexity, such as accounting for correlations. For instance, Denuit and Robert [9] explore the use of Ising models in collaborative insurance to capture such dependencies. Additionally, finding a way to update premiums dynamically is essential. This could be achieved either automatically—by scraping new data from competitors, insurance federations, or other relevant sources—or through decisions made by a "board" of customers and investors, who could leverage voting rights granted via tokens. While these refinements would increase costs, they seem practical given the potential savings on fraud and claim management.
The absence of staff turnover risk and the near-instantaneous resolution of claims offer significant advantages over traditional insurance models. High turnover in claims management teams, for instance, can lead to critical lapses in detecting changes in claim frequency.
In addition to systemic (market-wide) changes, insurers may encounter higher-than-expected claim frequencies. Adapting premiums using experience rating and credibility theory could help counteract potential adverse selection, provided the associated costs remain manageable. To complement this approach, the risk capital module should also be updated to reflect these credibility adjustments. This could involve leveraging large deviation principles rather than relying solely on central limit theorems, as discussed in Trufin and Loisel [25].
In both cases, it is worthwhile to establish a governance system where token-holding investors form a board with voting rights to decide on key parameters such as the safety loading level, premium principles, and the risk level $\alpha$. This aligns with the concept of decentralized organizations, which are extensively studied in management science and have naturally emerged within blockchain applications. However, classical issues such as over- or under-reactions stemming from individual risk perceptions and cognitive biases could be amplified, especially if voter participation is low or voting power is concentrated due to coalitions. These challenges are not uncommon, as low or uneven participation rates have frequently been observed in various decentralized networks, we refer the reader to the work of Messias et al. [20] for a comprehensive discussion.
As policyholders receive tokens when they underwrite a contract, they effectively become investors, creating a collaborative insurance framework. In this context, premium calculation could be based on the conditional mean risk-sharing principle, as proposed by Denuit and Robert [10].
Our smart contract is exposed to the risk of a "mass investor lapse," where investors may collectively withdraw their tokens. Stronger incentive mechanisms could be implemented to encourage investors to retain their tokens, particularly when the contract’s balance approaches the risk capital threshold. One potential approach is to introduce a coin age principle, where tokens held for longer periods gain additional value. This concept has already been explored in the implementation of the cryptocurrency PeerCoin [^11] See https://www.peercoin.net/. A traditional approach to managing the surplus of an insurance company is to implement a dividend policy that rewards investors or, more generally, token holders. This can be achieved using classical optimal dividend strategies. Barrier or band strategies, for instance, have been shown to be optimal across various modeling frameworks (see the survey by Avanzi [3]). In our framework, we operate within a discrete-time risk model with capital injections and withdrawals. This approach aligns with the structure of our system, where wealth is only inspected at discrete points—namely, during settlements or when an investor funds or withdraws from the contract.
In this paper, we have not addressed regulatory risks. In many countries, a smart contract like the one proposed would need to comply with insurance regulations. For instance, in the European Union, risk governance would need to be developed to align with the requirements of Pillar 2 of Solvency II, which presents significant challenges. Additionally, adhering to anti-money laundering (AML) regulations may necessitate incorporating a KYC (know your customer) module. While implementing such a module can be particularly challenging in a blockchain-based environment, some cryptocurrency wallet startups have successfully tackled this issue, providing potential avenues for further development.
In our approach, surplus management is handled through proxies referred to as the Minimum Capital Requirement (MCR) and Solvency Capital Requirement (SCR). If such a smart contract were to be classified as an insurance product, it would need to comply with the absolute minimum MCR threshold prescribed by Solvency II in Europe. This requirement could lead to a chicken-and-egg problem, as achieving the necessary critical mass of funds would be essential to meet the MCR. Additionally, calculating the SCR and MCR would typically involve integrating multiple risk modules, adding further complexity. Enhancing the simplified computations of these capital requirements would also incur additional implementation costs.
If the contract is not classified as an insurance product, it could lead to tax and regulatory implications. To ensure recognition as insurance, it is crucial to keep basis risk sufficiently low. Effective basis risk management not only protects policyholders but also mitigates the risk of the contract being reclassified as gambling. In some countries, policyholders may be required to provide brief proof of loss to validate claims. While we assume that the occurrence of a compensation-triggering event inherently results in a loss for the policyholder, real-world scenarios might necessitate additional efforts from either the policyholder or the smart contract’s governance system to address such requirements.
Sustainability considerations also warrant further investigation. While blockchain technology consumes significant amounts of electricity and water, its environmental impact could be partially offset by reducing the carbon emissions associated with traditional insurance operations, such as staffing, office infrastructure, and claim management processes.
In conclusion, we have developed a modest beta version of a smart contract for parametric insurance on the Ethereum blockchain. The Solidity code for the smart contract and the Python scripts for analyzing blockchain data are freely available in the smart__\__parametric__\__insurance GitHub repository. Beyond serving as a proof of concept, this work identifies several research questions that need to be addressed to make the product fully operational. We hope that our open-source beta version will provide a foundation for other researchers to build upon, facilitating the development of enhanced features and overcoming the limitations we have outlined.
Acknowledgements
Pierre-O’s work is conducted as part of the Research Chair DIALOG [^12] See https://chaire-dialog.fr/en/welcome/, under the auspices of the Risk Foundation, an initiative by CNP Assurances. His research is also supported by the ANR project BLOCKFI [^13] See https://anr.fr/Project-ANR-24-CE38-7885. Stéphane acknowledges support from the Research Chair ACTIONS [^14] See https://chaireactions.fr/ (funded by BNP Paribas Cardif), the Initiative de Recherche Sustainable Actuarial Science and Climatic Risk [^15] See https://sites.google.com/view/stephaneloisel/recherche/projets/actuariat-durable-et-risques-climatiques (funded by Milliman Paris), and the ANR project DREAMES [^16] See https://anr.fr/Projet-ANR-21-CE46-0002.